OpenStudy (anonymous):

composite function w/ sqrt: 4(sqrt(2x-1)^2+3(sqrt(2x-1))+5 I can clear the first sqrt but not the second; please help.

OpenStudy (anonymous):

\[h(x)= 4(\sqrt{2x-1})^2+3(\sqrt{2x-1})+5\] \[g(x)=4x^2+3x+5\] \[f(x)=\sqrt{2x-1}\] \[h=gof\]

OpenStudy (anonymous):

is this what you are asking

OpenStudy (anonymous):

Unfortunately no. I need the new composite function; fog. The answer is supposed to be cleared of all sqrt's. My answer has been 64x^2-18x+10. I got there by flipping poly's across the equals sign, squaring each side and moving terms again to derive an answer. I need verification that this is the correct procedure.

OpenStudy (anonymous):

what is the function you gave on top g(x) f(x) gof(x) or what

OpenStudy (anonymous):

(fog)(x); f(x)=4x^2-3x+5 g(x)= sqrt(2x-1)

OpenStudy (anonymous):

\[4(2x-1)-3(\sqrt{2x-1})+5\] the 2nd square rot does not have to go away

OpenStudy (anonymous):

That's what I thought, but my instructor says it does.

OpenStudy (anonymous):

\[8x-3(2x-1)^{1/2}+1\] thats the only thing i can do

OpenStudy (anonymous):

Thanks. The program rejects that as the correct answer. I'm stumped, but I appreciate your help. By the way, what program are you using to type your math equations?